OpenStudy (anonymous):

solve the absolute value equation (show steps) |3k-3|+8=15

OpenStudy (anonymous):

-7=3k-3=7 -4=3k=10 -4/3=k=10/3

OpenStudy (anonymous):

great

OpenStudy (anonymous):

|3k-3| = 7 (3k-3) = 7 or (-(3k-3)) = 7 3k = 10 or -3k = 4 k = 10/3 or k = -4/3

OpenStudy (anonymous):

^^ what are you talking about look at my solution -4/3=k=10/3

OpenStudy (anonymous):

Bnut.. k cannot equal -4/3 AND 10/3. It must be one or the other. Either one is a solution, but they are not equal.

OpenStudy (anonymous):

its the way i wrote it, instead of using two lines I solved it in one step instead of two thats all polpak

OpenStudy (anonymous):

I understood what you meant, but you wrote -4/3 = k = 10/3. And those are definitely not equal.

OpenStudy (anonymous):

well obviously haha im sure the person i helped out understood or else he would have said something thats what im here to do, not to punctuate people, its to make thing CLEAR so people can have help not to talk about here-say or mislead people

OpenStudy (anonymous):

You can say something like: \(3k -3 \in \{-7,7\}\)\[\implies 3k \in \{-4,10\}\]\[\implies k \in \{-\frac{4}{3}, \frac{10}{3}\}\] But saying that those are all equal is confusing.

OpenStudy (anonymous):

ok thank